# Intercept Cancer: Cancer Pre-Screening with Large Scale Healthcare Foundation Models

### v1

@@ Appendix C EHRSHOT
Dataset and Prediction Task Overview. EHRSHOT is the latest published EHR dataset for evaluating the few-shot performance of foundation models on clinical prediction tasks. It consists of structured, de-identified, and longitudinal Electronic Health Records (EHRs) from 6,739 patients treated at Stanford Medicine, including diagnosis codes, procedures, medications, and laboratory test results. EHRSHOT is sourced from the EHR system in both Stanford Health Care and Lucile Packard Children’s Hospital, where 2.57M patients used to pretrain CLMBR, a clinical foundation model used to evaluate the EHRSHOT benchmark in the original paper [97]. It contains 15 distinct clinical prediction tasks under the four main categories:
@@
where the one-year pancreatic cancer prediction is under the category "Assignment of New Diagnoses". It is worth noting that, pancreatic cancer prediction is the only available cancer prediction task in EHRSHOT. Each designed to test the ability of foundation models for accurate predictions under limited labeled data at the time of patient visit. Refer Table 3 in the original EHRSHOT paper [97] for complete task description and statistics.
@@
Mapping EHRSHOT to NHIRD. ICD-based codes such as ICD-9-Proc (ICD-9 Procedure code), ICD-10-PCS (ICD-10 Procedure code), and ICD-O-3 (ICD oncology code) can be mapped to ICD-9 using publicly available mapping from the Centers for Medicare & Medicaid Services (CMS). Diagnostic concepts in SNOMED can also be mapped to ICD-9 via the "ICD-9-CM Diagnostic Codes to SNOMED CT Map" provided by the U.S. National Library of Medicine (NLM). However, other coding systems, e.g., SNOMED procedures, CPT (coding for medical services, procedures, and other practices), and RxNorm (codes for drug prescriptions), lack direct mappings to the NHIRD coding scheme. To address this, we adopt a text-based semantic matching approach by embedding the textual descriptions of codes using Sentence Transformers [84] and formulating the mapping as a dense retrieval problem, where several prior works adopt this method to align different medical concepts and terms [81; 56; 104]. Codes from each coding system are matched to the most semantically similar code in NHIRD based on text similarity. We use Faiss as our dense retrieval framework [22]. Table 12 summarizes the mapping strategies and source-target relationships between each coding system. With the fundamental structural and terminological differences between these two EHR systems, our average top-1 semantic alignment score (cosine similarity) between code in EHRSHOT and code in NHIRD is approximately 84.3%. This score reflects the inherent challenge of mapping between heterogeneous and non-standardized EHR codebases, especially when deploying healthcare foundation models across healthcare systems and populations, showing the difficulties of model generalization.
Appendix D NPV evaluation
Here we present the formula to calculate Negative Predictive Value (NPV) given screening model Sensitivity, Specificity, and Prevalence [59]:
where Sensitivity, Specificity, and Prevalence are in decimal, and we use positive rate for each target cancer as Prevalence.
@@ Appendix D NPV evaluation
Appendix E Implementation Details
@@
Baselines. For the tree-based baselines, XGBoost and LightGBM, we utilize their official Python packages, xgboost 111https://xgboost.readthedocs.io/en/stable/index.html and lightgbm 222https://lightgbm.readthedocs.io/en/stable/, respectively. For deep learning baselines, we employ the PyHealth framework [101] and perform supervised learning tasks directly on the cancer screening benchmarks. For language model baselines, we convert medical codes into text using language model vocabularies and concatenate them into sequences representing patient histories for fine-tuning. In NHIRD, ICD9 codes are mapped to standardized textual descriptions using official code mapping tables. Each medical code, representing a diagnosis, surgery, treatment, or medication, is transformed accordingly. Patient histories are then constructed by concatenating the medical text, similar to processing text documents in language models. The sequence length limit is 1024 for BioGPT and 2048 for Qwen.
@@
| Cancer: Pancreatic, Positive/Negative (Incidence Ratio): 1201/30382 (3.95%) | |||||
| 0.999 | 0.001 | 0.600 | 0.999 | 0.956 | 25.146 |
| 0.999 | 0.001 | 0.630 | 0.999 | 0.945 | 24.853 |
| 0.997 | 0.002 | 0.654 | 0.998 | 0.932 | 24.517 |
| 0.990 | 0.003 | 0.670 | 0.997 | 0.909 | 23.893 |
| 0.975 | 0.003 | 0.684 | 0.997 | 0.886 | 23.293 |
| 0.941 | 0.004 | 0.696 | 0.996 | 0.868 | 22.826 |
| 0.825 | 0.005 | 0.707 | 0.995 | 0.836 | 21.975 |
| 0.710 | 0.007 | 0.717 | 0.993 | 0.806 | 21.220 |
| 0.550 | 0.009 | 0.728 | 0.991 | 0.780 | 20.521 |
| 0.391 | 0.010 | 0.740 | 0.990 | 0.753 | 19.796 |
| Cancer: Liver, Positive/Negative (Incidence Ratio): 1099/28353 (3.88%) | |||||
| 0.915 | 0.004 | 0.601 | 0.996 | 0.856 | 22.941 |
| 0.891 | 0.004 | 0.612 | 0.996 | 0.848 | 22.715 |
| 0.852 | 0.004 | 0.624 | 0.996 | 0.833 | 22.338 |
| 0.819 | 0.005 | 0.637 | 0.994 | 0.823 | 22.070 |
| 0.757 | 0.006 | 0.646 | 0.994 | 0.807 | 21.622 |
| 0.705 | 0.006 | 0.654 | 0.993 | 0.791 | 21.197 |
| 0.652 | 0.007 | 0.668 | 0.993 | 0.782 | 20.949 |
| 0.617 | 0.008 | 0.676 | 0.992 | 0.765 | 20.479 |
| 0.579 | 0.009 | 0.684 | 0.991 | 0.754 | 20.213 |
| 0.508 | 0.010 | 0.696 | 0.990 | 0.736 | 19.713 |
| Cancer: Lung, Positive/Negative (Incidence Ratio): 868/30153 (2.88%) | |||||
| 0.853 | 0.006 | 0.600 | 0.994 | 0.730 | 26.078 |
| 0.818 | 0.007 | 0.604 | 0.993 | 0.725 | 25.902 |
| 0.727 | 0.007 | 0.608 | 0.992 | 0.713 | 25.465 |
| 0.686 | 0.007 | 0.614 | 0.993 | 0.705 | 25.196 |
| 0.632 | 0.008 | 0.619 | 0.992 | 0.698 | 24.957 |
| 0.572 | 0.008 | 0.622 | 0.992 | 0.687 | 24.522 |
| 0.523 | 0.008 | 0.628 | 0.992 | 0.680 | 24.316 |
| 0.437 | 0.009 | 0.630 | 0.991 | 0.663 | 23.696 |
| 0.395 | 0.010 | 0.634 | 0.990 | 0.657 | 23.484 |
| 0.335 | 0.010 | 0.637 | 0.990 | 0.648 | 23.169 |
@@
| Cancer: Pancreatic, Positive/Negative (Incidence Ratio): 1201/30382 (3.95%) | ||||||
| CATCH-FM-1b (Code Sequence) | 0.391 | 0.01 | 0.740 | 0.99 | 0.753 | 19.8 |
| XGBoost (Bag-of-words) | 0.327 | 0.01 | 0.572 | 0.99 | 0.695 | 18.3 |
| Qwen (Language Sequence) | 0.320 | 0.01 | 0.627 | 0.99 | 0.716 | 18.8 |
| Cancer: Liver, Positive/Negative (Incidence Ratio): 1099/28353 (3.88%) | ||||||
| CATCH-FM-1b (Code Sequence) | 0.508 | 0.01 | 0.696 | 0.99 | 0.736 | 19.7 |
| XGBoost (Bag-of-words) | 0.297 | 0.01 | 0.614 | 0.99 | 0.708 | 19.0 |
| Qwen (Language Sequence) | 0.239 | 0.01 | 0.591 | 0.99 | 0.696 | 18.7 |
| Cancer: Lung, Positive/Negative (Incidence Ratio): 868/30153 (2.88%) | ||||||
| CATCH-FM-1b (Code Sequence) | 0.335 | 0.01 | 0.637 | 0.99 | 0.648 | 23.2 |
| XGBoost (Bag-of-words) | 0.258 | 0.01 | 0.586 | 0.99 | 0.628 | 22.4 |
| Qwen (Language Sequence) | 0.558 | 0.01 | 0.621 | 0.99 | 0.642 | 22.9 |
Appendix F Analysis of Threshold Tuning
@@
Clinically, models often adjust decision thresholds to prioritize sensitivity or specificity [18; 76], such as reducing false positives in prescreening or increasing sensitivity in high-risk groups. Figure 8 shows that CATCH-FM supports tunable sensitivity and specificity, which makes it well-suited for diverse clinical settings.
@@
We evaluate our model and carefully select decision thresholds from the AUROC curve to ensure a true positive rate (TPR) 0.6, a false positive rate (FPR) 0.01, and a specificity of at least 0.99. We also report Relative Risk (RR), following suggested work in Placido et al. [65], to quantify the odds of cancer among those classified as positive compared to a random pick based just on the population disease incidence. We provide a complete threshold selection results in Table 16.
@@
Under a fixed specificity of 0.99 and FPR of 0.01, our model demonstrates strong performance across all three cancer prediction tasks. For pancreatic cancer, a threshold of 0.39 yields a TPR of 0.74, and a relative risk of 19.8. For liver cancer, a threshold of 0.51 yields a TPR of 0.696 and a relative risk of 19.7. For lung cancer, a threshold of 0.34 yields a TPR of 0.64 and a relative risk of 23.2. These results highlight the utility of our method across diverse cancer types. Under the fixed specificity threshold, we ensure a fair comparison with the baselines, as detailed in Table 17. Across all metrics, our model consistently outperforms the baselines.
@@
To simulate clinical implementation, following the approach proposed in Placido et al. [65], we adopt an operational decision point that simulates cost constraints, where only the top 0.1% of patients (by predicted risk) are eligible to be advanced to a surveillance program. At this threshold, our model achieves relative risk scores of 24.6, 24.9, and 35.74 for pancreatic, liver, and lung cancers prediction, respectively. This also further indicates the clinical utility of our model. We also report AUROC, AUPRC, and Relative Risk curves across thresholds to compare our model with baseline methods (XGBoost with bag-of-words and Qwen with language sequence input), as shown in Figure 9.
@@
Appendix G Benefits of Compute-Optimal Pretraining
@@
Appendix H Analysis of Cohort Control
@@
Appendix I Time Window Exclusion
@@
| Category | Pos./Neg. (ratio) | F1 | AUROC | AUPRC | Specificity | Sensitivity |
| Cancer: Pancreatic | ||||||
| Demographics: Gender | ||||||
| Male | 616/17684 (3.48%) | 86.3 | 97.0 | 80.1 | 99.2 | 71.2 |
| Female | 584/12570 (4.65%) | 88.5 | 97.1 | 85.3 | 99.1 | 75.9 |
| Demographics: Age | ||||||
| 20-40 | 136/11018 (1.23%) | 85.9 | 93.9 | 73.8 | 99.7 | 67.8 |
| 40-60 | 434/4975 (8.72%) | 86.7 | 94.0 | 81.7 | 98.3 | 72.1 |
| 60-80 | 540/2248 (24.02%) | 86.4 | 93.7 | 87.4 | 94.8 | 77.8 |
| 80+ | 79/246 (32.11%) | 81.2 | 87.4 | 81.5 | 91.8 | 69.4 |
| Clinical Covariates: Visit Frequency | ||||||
| 100 | 627/27489 (2.28%) | 82.4 | 95.7 | 70.5 | 99.4 | 61.1 |
| 100-200 | 366/2403 (15.23%) | 90.7 | 97.6 | 92.3 | 97.4 | 84.8 |
| 200-300 | 132/374 (35.29%) | 92.7 | 98.6 | 96.8 | 95.2 | 91.4 |
| 300+ | 76/116 (65.52%) | 92.0 | 98.3 | 97.5 | 93.3 | 90.8 |
| Cancer: Liver | ||||||
| Demographics: Gender | ||||||
| Male | 690/16631 (4.15%) | 85.0 | 94.9 | 76.5 | 99.1 | 67.2 |
| Female | 408/11622 (3.51%) | 85.2 | 95.5 | 78.1 | 99.0 | 70.9 |
| Demographics: Age | ||||||
| 20-40 | 163/10072 (1.62%) | 73.4 | 87.9 | 48.2 | 99.5 | 41.3 |
| 40-60 | 441/4728 (9.33%) | 84.5 | 92.8 | 78.7 | 97.6 | 70.0 |
| 60-80 | 450/2072 (21.72%) | 88.2 | 94.2 | 88.1 | 96.1 | 79.4 |
| 80+ | 30/223 (13.45%) | 73.6 | 77.4 | 58.7 | 93.2 | 55.3 |
| Clinical Covariates: Visit Frequency | ||||||
| 100 | 566/25828 (2.19%) | 78.6 | 91.6 | 60.6 | 99.4 | 52.0 |
| 100-200 | 316/2070 (15.27%) | 89.1 | 97.7 | 90.4 | 96.7 | 83.3 |
| 200-300 | 134/360 (37.22%) | 90.0 | 97.8 | 95.0 | 91.8 | 91.8 |
| 300+ | 83/95 (87.37%) | 88.4 | 94.8 | 94.7 | 89.5 | 87.2 |
| Cancer: Lung | ||||||
| Demographics: Gender | ||||||
| Male | 498/16457 (3.03%) | 82.4 | 92.9 | 68.7 | 99.2 | 61.4 |
| Female | 369/13622 (2.71%) | 82.6 | 92.8 | 69.9 | 99.5 | 60.0 |
| Demographics: Age | ||||||
| 20-40 | 80/11654 (0.69%) | 72.7 | 78.4 | 36.2 | 99.9 | 32.3 |
| 40-60 | 293/5680 (6.16%) | 80.0 | 89.3 | 65.6 | 98.9 | 54.1 |
| 60-80 | 420/2324 (18.07%) | 82.9 | 90.5 | 78.6 | 94.8 | 70.8 |
| 80+ | 68/231 (29.44%) | 79.4 | 86.1 | 77.2 | 92.2 | 65.0 |
| Clinical Covariates: Visit Frequency | ||||||
| 100 | 401/27458 (1.46%) | 74.0 | 87.8 | 46.9 | 99.6 | 40.6 |
| 100-200 | 267/2295 (11.63%) | 86.4 | 94.8 | 83.6 | 97.2 | 75.5 |
| 200-300 | 127/314 (40.45%) | 87.8 | 95.5 | 92.4 | 91.9 | 84.7 |
| 300+ | 73/86 (84.88%) | 80.7 | 89.9 | 88.8 | 85.6 | 75.6 |
Appendix J Analysis of Demographic Subgroups
Table 21 summarizes the performance of CATCH-FM-1b across demographic subgroups. No significant bias is observed in terms of gender, as performance metrics (F1, AUROC, AUPRC) are similar between males and females across all cancers. Model performance peaks in middle-aged groups (40-60 and 60-80), with reduced sensitivity and F1 scores in younger (20-40) and older (80+) groups, likely due to these groups presenting less distinct or atypical clinical features, making detection more challenging. For the younger cohort, the limited number of positive cases may lead to data scarcity issues, resulting in reduced model sensitivity or generalizability. In contrast, the older group may exhibit more clinical complexity and comorbidities, making accurate prediction more challenging despite a higher prevalence rate. A more deep investigation is an interesting direction for future research.
| Model | FLOPs | F1 (Macro) | Accuracy | AUC | AUPRC | Specificity | Sensitivity* |
| Disease: Heart Attack, Positive/Negative (Ratio): 10205/429248 (2.38%) | |||||||
| XGBoost | - | 0.734 | 0.982 | 0.953 | 0.554 | 0.997 | 0.357 |
| LightGBM | - | 0.742 | 0.982 | 0.957 | 0.575 | 0.997 | 0.375 |
| CATCH-FM-160m | 1e18 | 0.771 | 0.981 | 0.951 | 0.579 | 0.992 | 0.508 |
| CATCH-FM-410m | 2e19 | 0.775 | 0.98 | 0.935 | 0.558 | 0.991 | 0.516 |
| CATCH-FM-720m | 4e19 | 0.777 | 0.981 | 0.935 | 0.576 | 0.992 | 0.522 |
| CATCH-FM-1b | 2e20 | 0.775 | 0.98 | 0.95 | 0.58 | 0.991 | 0.538 |
| Disease: Stroke, Positive/Negative (Ratio): 21815/625873 (3.49%) | |||||||
| XGBoost | - | 0.84 | 0.982 | 0.975 | 0.765 | 0.995 | 0.599 |
| LightGBM | - | 0.854 | 0.983 | 0.977 | 0.788 | 0.995 | 0.646 |
| CATCH-FM-160m | 1e18 | 0.873 | 0.984 | 0.982 | 0.827 | 0.994 | 0.714 |
| CATCH-FM-410m | 2e19 | 0.873 | 0.984 | 0.982 | 0.827 | 0.994 | 0.714 |
| CATCH-FM-720m | 4e19 | 0.878 | 0.985 | 0.983 | 0.834 | 0.994 | 0.727 |
| CATCH-FM-1b | 2e20 | 0.874 | 0.984 | 0.974 | 0.785 | 0.992 | 0.755 |
Appendix K Performance of Heart Attack and Stroke Screening
Building on CATCH-FM’s success in cancer screening, we extend its evaluation to other life-threatening conditions—heart attack and stroke, both leading global causes of death [60] and prevalent research topics in NHIRD [38; 10; 52; 9; 88; 86]. Early screening tools for these conditions are crucial for timely interventions. We perform binary classification to predict whether a patient will experience a stroke or heart attack exactly one year after their current clinical visit, using the preceding five years of medical history. To ensure predictions emphasize early screening efficacy, the target event (1 for detected, 0 for not detected) is strictly set at the one-year mark. Following prior work, we use ICD9 codes 43301, 43311, 43321, 43331, 43381, 43391, 43401, 43411, and 43491 for stroke [38] and the first three digits 410 for heart attack [10], applying the same case-control matching criteria as in Section 6.1. Consistent with our cancer screening approach, we evaluate CATCH-FM at compute-optimal scales and compare it against tree-based baselines.
As shown in Table 22, CATCH-FM demonstrates superior performance and outperforms all baselines with high sensitivity at 99% specificity, providing reliable early warnings for both life-threatening medical conditions. Notably, the scaling behaviors on these tasks mirror the observed scaling laws in cancer screening, further underscoring CATCH-FM’s robustness and adaptability. These findings highlight its potential for deployment in diverse, life-saving clinical scenarios. They also emphasize the value of scaling larger models with expanded medical datasets to create reliable healthcare foundation models that address critical medical use cases.
Appendix L Details of Interpretability Experiments
We process positive patients’ event token sequences using the fine-tuned CATCH-FM-1b model on each cancer dataset to obtain hidden states for every token. Following [43], we then train a TopK sparse autoencoder (SAE) on the hidden states of the [EOS] token, , which serves as an aggregated representation of patient trajectories. The SAE is implemented as follows:

### v2

@@ Appendix C EHRSHOT
Dataset and Prediction Task Overview. EHRSHOT is the latest published EHR dataset for evaluating the few-shot performance of foundation models on clinical prediction tasks. It consists of structured, de-identified, and longitudinal Electronic Health Records (EHRs) from 6,739 patients treated at Stanford Medicine, including diagnosis codes, procedures, medications, and laboratory test results. EHRSHOT is sourced from the EHR system in both Stanford Health Care and Lucile Packard Children’s Hospital, where 2.57M patients used to pretrain CLMBR, a clinical foundation model used to evaluate the EHRSHOT benchmark in the original paper (Wornow et al., 2023). It contains 15 distinct clinical prediction tasks under the four main categories:
@@
where the one-year pancreatic cancer prediction is under the category ”Assignment of New Diagnoses”. It is worth noting that, pancreatic cancer prediction is the only available cancer prediction task in EHRSHOT. Each designed to test the ability of foundation models for accurate predictions under limited labeled data at the time of patient visit. Refer Table 3 in the original EHRSHOT paper (Wornow et al., 2023) for complete task description and statistics.
@@
Mapping EHRSHOT to NHIRD. As CATCH-FM is pretrained on NHIRD using ICD-9 and NHI codes, we map SNOMED to ICD-9 using official CMS and NLM mappings. While 43% of codes have one-to-one mappings (mostly SNOMED to ICD-9), the remaining 57% (e.g., RxNorm, CPT) are aligned to NHI codes via semantic matching on text descriptions (Liu et al., 2020; Sung et al., 2020a; Yuan et al., 2022). Due to differences in drug, procedure, and medical codes, the average cosine similarity between EHRSHOT and top-1 matched NHIRD codes is 84.3%, posing a significant generalization challenge where no gold-mapping standard exists.
ICD-based codes such as ICD-9-Proc (ICD-9 Procedure code), ICD-10-PCS (ICD-10 Procedure code), and ICD-O-3 (ICD oncology code) can be mapped to ICD-9 using publicly available mappings from the Centers for Medicare & Medicaid Services (CMS). Diagnostic concepts in SNOMED can also be mapped to ICD-9 via the “ICD-9-CM Diagnostic Codes to SNOMED CT Map” provided by the U.S. National Library of Medicine (NLM). However, other coding systems, e.g., SNOMED procedures, CPT (coding for medical services, procedures, and other practices), and RxNorm (codes for drug prescriptions), lack direct mappings to the NHIRD coding scheme. Table 13 summarizes the mapping strategies and source-target relationships between each coding system. To address this, we adopt a text-based semantic matching approach by embedding111We adopt Salesforce/SFR-Embedding-Mistral as encoder from Huggingface. the textual descriptions of codes using Sentence Transformers (Thakur et al., 2021) and formulating the mapping as a dense retrieval problem, a method widely used to align medical concepts across ontologies (Sung et al., 2020a; Liu et al., 2020; Yuan et al., 2022). Codes from each system are matched to the most semantically similar code in NHIRD based on text similarity, using Faiss as the retrieval backend (Douze et al., 2024).
| Code in EHRSHOT | Mapping methods | Code in NHIRD |
| SNOMED | SNOMED to ICD9 mapping | ICD-9-CM |
| SNOMED - physical object | Semantic text matching | NHI Order Code |
| ICDO3 | ICD mapping by CMS | ICD-9-CM |
| ICD10-Proc | ICD mapping by CMS | ICD-9-Proc |
| RxNorm | Semantic text matching | NHI Drug Code |
| CPT | Semantic text matching | NHI Order / Service Code |
| HCPCS | Semantic text matching | NHI Material Code |
| LONIC | Semantic text matching | NHI Order Code |
| Code type (Total) | Exact | 0.98 threshold |
| SNOMED (11598) | 8998 (77.6%) | 9072 (78.2%) |
| ICD10PCS (3669) | 3618 (98.6%) | 3618 (98.6%) |
| ICD03 (96) | 76 (79.2%) | 83 (86.5%) |
| RxNorm (5433) | 0 (0.0%) | 0 (0.0%) |
| CPT (4675) | 0 (0.0%) | 3 (0.06%) |
| LONIC (3945) | 2 (0.05%) | 3 (0.08%) |
| HCPCS (64) | 0 (0.0%) | 0 (0.0%) |
| All (29480) | 12694 (43.1%) | 12779 (43.4%) |
Since no gold mapping standard exists for semantic alignment, it is difficult to determine whether moderate similarity scores (e.g., 0.7–0.85) reflect true mappings or noise. To approximate exact mappings and ensure high precision, we therefore adopt a strict cutoff of 0.98 as the threshold for soft matching. Table 14 summarizes the resulting coverage under different matching strategies. Even with semantic matching, the final coverage reaches only 43%, implying that over half of the medical information is lost. This substantial loss reflects the inherent challenge of aligning heterogeneous and non-standardized EHR format, especially when deploying healthcare foundation models across different healthcare systems and populations, and underscores the difficulties of achieving robust model generalization.
@@
Appendix D Implementation Details
@@
Baselines. For the tree-based baselines, XGBoost and LightGBM, we utilize their official Python packages, xgboost 222https://xgboost.readthedocs.io/en/stable/index.html and lightgbm 333https://lightgbm.readthedocs.io/en/stable/, respectively. For deep learning baselines, we employ the PyHealth framework (Yang et al., 2023a) and perform supervised learning tasks directly on the cancer screening benchmarks. For language model baselines, we convert medical codes into text using language model vocabularies and concatenate them into sequences representing patient histories for fine-tuning. In NHIRD, ICD9 codes are mapped to standardized textual descriptions using official code mapping tables. Each medical code, representing a diagnosis, surgery, treatment, or medication, is transformed accordingly. Patient histories are then constructed by concatenating the medical text, similar to processing text documents in language models. The sequence length limit is 1024 for BioGPT and 2048 for Qwen.
@@
| Cancer: Pancreatic, Positive/Negative (Incidence Ratio): 452/28058 (1.59%) | |||||
| 0.996 | 0.000 | 0.188 | 1.000 | 1.000 | 63.075 |
| 0.980 | 0.000 | 0.330 | 0.9999 | 0.974 | 61.426 |
| 0.932 | 0.001 | 0.407 | 0.9994 | 0.911 | 57.455 |
| 0.815 | 0.001 | 0.458 | 0.9986 | 0.841 | 53.075 |
| 0.686 | 0.002 | 0.491 | 0.9975 | 0.760 | 47.954 |
| 0.495 | 0.004 | 0.535 | 0.9960 | 0.684 | 43.119 |
| 0.411 | 0.005 | 0.566 | 0.9946 | 0.627 | 39.577 |
| 0.264 | 0.008 | 0.586 | 0.9918 | 0.536 | 33.836 |
| 0.204 | 0.010 | 0.606 | 0.9900 | 0.495 | 31.196 |
| Cancer: Liver, Positive/Negative (Incidence Ratio): 509/26148 (1.95%) | |||||
| 0.984 | 0.000 | 0.134 | 0.9999 | 0.958 | 50.158 |
| 0.957 | 0.001 | 0.242 | 0.9994 | 0.891 | 46.679 |
| 0.905 | 0.001 | 0.330 | 0.9985 | 0.816 | 42.711 |
| 0.841 | 0.003 | 0.381 | 0.9972 | 0.727 | 38.053 |
| 0.768 | 0.004 | 0.422 | 0.9958 | 0.660 | 34.539 |
| 0.702 | 0.006 | 0.446 | 0.9945 | 0.612 | 32.044 |
| 0.635 | 0.007 | 0.477 | 0.9932 | 0.576 | 30.157 |
| 0.588 | 0.008 | 0.501 | 0.9919 | 0.545 | 28.536 |
| 0.515 | 0.010 | 0.536 | 0.9901 | 0.512 | 26.824 |
| Cancer: Lung, Positive/Negative (Incidence Ratio): 868/30153 (1.45%) | |||||
| 0.997 | 0.000 | 0.130 | 0.9998 | 0.897 | 62.455 |
| 0.992 | 0.001 | 0.243 | 0.9995 | 0.866 | 60.332 |
| 0.977 | 0.001 | 0.336 | 0.9992 | 0.854 | 59.456 |
| 0.950 | 0.002 | 0.366 | 0.9982 | 0.749 | 52.157 |
| 0.914 | 0.003 | 0.414 | 0.9974 | 0.702 | 48.911 |
| 0.825 | 0.004 | 0.454 | 0.9962 | 0.633 | 44.087 |
| 0.716 | 0.005 | 0.491 | 0.9950 | 0.589 | 41.002 |
| 0.593 | 0.007 | 0.509 | 0.9931 | 0.519 | 36.167 |
| 0.385 | 0.010 | 0.531 | 0.9901 | 0.439 | 30.576 |
@@
| Cancer: Pancreatic, Positive/Negative (Incidence Ratio): 452/28058 (1.59%) | ||||||
| CATCH-FM-2.4b (Code Sequence) | 0.204 | 0.01 | 0.606 | 0.99 | 0.495 | 31.196 |
| XGBoost (Bag-of-words) | 0.180 | 0.01 | 0.310 | 0.99 | 0.335 | 21.126 |
| Qwen (Language Sequence) | 0.408 | 0.01 | 0.254 | 0.99 | 0.324 | 17.644 |
| Cancer: Liver, Positive/Negative (Incidence Ratio): 509/26148 (1.95%) | ||||||
| CATCH-FM-2.4b (Code Sequence) | 0.515 | 0.01 | 0.536 | 0.99 | 0.512 | 26.824 |
| XGBoost (Bag-of-words) | 0.200 | 0.01 | 0.363 | 0.99 | 0.415 | 21.723 |
| Qwen (Language Sequence) | 0.437 | 0.01 | 0.324 | 0.99 | 0.421 | 19.3 |
| Cancer: Lung, Positive/Negative (Incidence Ratio): 868/30153 (1.45%) | ||||||
| CATCH-FM-2.4b (Code Sequence) | 0.385 | 0.01 | 0.531 | 0.99 | 0.439 | 30.576 |
| XGBoost (Bag-of-words) | 0.172 | 0.01 | 0.323 | 0.99 | 0.323 | 22.5 |
| Qwen (Language Sequence) | 0.573 | 0.01 | 0.188 | 0.99 | 0.238 | 15.239 |
Appendix E Analysis of Threshold Tuning
@@
Clinically, models often adjust decision thresholds to prioritize sensitivity or specificity (Collins & Moons, 2012; Steyerberg & Steyerberg, 2019), such as reducing false positives in prescreening or increasing sensitivity in high-risk groups. Figure 8 shows that CATCH-FM supports tunable sensitivity and specificity, which makes it well-suited for diverse clinical settings. It show that CATCH-FM maintains its advantage across all thresholds, confirming its ability to identify high-risk patients for further screening (high sensitivity) and avoid unnecessary patient distress (high specificity). We also report AUROC curves across thresholds to compare our model with baseline methods (XGBoost and Qwen), as shown in Figure 9.
@@
We evaluate our model and carefully select decision thresholds from the AUROC curve to ensure a specificity of at least 0.99. We also report Relative Risk (RR), to quantify the odds of cancer among those classified as positive compared to a random pick based just on the population disease incidence following suggested work in Placido et al. (2023a). We provide a complete threshold selection results in Table 18.
@@
Under a fixed specificity of 0.99 and FPR of 0.01, our model demonstrates strong performance across all three cancer prediction tasks. For pancreatic cancer, a threshold of 0.20 yields a TPR of 0.61, and a relative risk of 31.2. For liver cancer, a threshold of 0.52 yields a TPR of 0.54 and a relative risk of 26.8. For lung cancer, a threshold of 0.39 yields a TPR of 0.53 and a relative risk of 30.6. These results highlight the utility of our method across diverse cancer types. Under the fixed specificity threshold, we ensure a fair comparison with the baselines, as detailed in Table 19. Across all metrics, our model consistently outperforms the baselines.
@@
To simulate clinical implementation, following the approach proposed in Placido et al. (2023a), we adopt an operational decision point that simulates cost constraints, where only the top 0.1% of patients (by predicted risk) are eligible to be advanced to a surveillance program. At this threshold, our model achieves relative risk scores of 63.1, 52.0, and 61.9 for pancreatic, liver, and lung cancers prediction, respectively. This also further indicates the clinical utility of our model.
@@
Appendix F Benefits of Compute-Optimal Pretraining
@@
Appendix G Analysis of Cohort Control
@@
Appendix H Time Window Exclusion
@@
Appendix I Details of Interpretability Experiments
We process positive patients’ event token sequences using the fine-tuned CATCH-FM-1b model on each cancer dataset to obtain hidden states for every token. Following (Kang et al., 2024), we then train a TopK sparse autoencoder (SAE) on the hidden states of the [EOS] token, , which serves as an aggregated representation of patient trajectories. The SAE is implemented as follows: